### v2

# Searching for dark matter signals with high energy astrophysical neutrinos in IceCube

Abstract
High-energy neutrinos provide a potentially powerful and distinctive probe for dark matter (DM) - neutrino interactions, particularly in environments with enhanced DM densities, such as the DM spikes predicted to form around supermassive black holes (SMBHs) at the center of active galactic nuclei (AGN). Recent observations by the IceCube Neutrino Observatory, which identified four AGN, namely TXS 0506+056, NGC 1068, PKS 1424+240, and NGC 4151 as neutrino sources, provide a unique opportunity to search for signatures of these interactions. In this study, we use IceCube data to derive the most stringent constraints to date on both the energy-dependent and energy-independent DM-neutrino scattering cross-sections. We perform a statistical analysis using data from individual sources as well as a combined (stacked) analysis of all four sources. Our strongest limits arise from the stacking analysis, yielding an upper bound of $\sigma_{0}\lesssim 8\times 10^{-39}$ cm2 for an energy-independent cross-section and $\sigma_{0}\lesssim 10^{-39}$ cm2 for a linearly energy-dependent cross-section, both at 90% confidence level, particularly in scenarios involving the adiabatic growth of black holes.
preprint: APS/123-QED
I Introduction
The existence of dark matter (DM) is well-established through its gravitational interactions with standard model (SM) matter in astrophysical and cosmological systems [1, 2]. Yet, its fundamental nature remains unknown. Discovery of a clear signal necessitates a multi-pronged approach that integrates novel theoretical techniques with a broad experimental search program. While both neutrinos and DM may interact only feebly with the rest of the SM, it is reasonable to explore the possibility that they interact strongly with each other. Among the many probes, high-energy astrophysical neutrinos present a remarkable opportunity to investigate possible strong interactions between these two sectors, across vast astrophysical distances [3, 4, 5, 6, 7, 8]. A fraction of these high-energy neutrinos are created in the extreme and dense environments of AGN, which harbor supermassive black holes (SMBHs), and where conditions are ripe for acceleration and interactions of cosmic-ray protons and nuclei [9, 10, 11, 12, 13, 14, 15]. As these neutrinos propagate through space, interactions with DM, in regions of high DM density, may lead to observable scattering or attenuation effects, particularly in the intriguing and predicted DM density spikes surrounding SMBHs [6, 7].
Recently, the IceCube Neutrino Observatory has made groundbreaking strides in neutrino astronomy by detecting several extragalactic point sources, including the notable TXS 0506+056 [16, 17], NGC 1068 [18], PKS 1424+240 [18], and NGC 4151 [19]. These findings represent a major advancement, allowing for further direct tests of DM-neutrino scattering. In this study, we leverage publicly available IceCube data to derive constraints on the DM-neutrino scattering cross-section, considering both energy-independent and energy-dependent interaction scenarios. To this end, we analyze individual sources and perform a stacking analysis that integrates their results to obtain comprehensive bounds on DM-neutrino scattering.
Previous studies have used astrophysical neutrino data from TXS 0506+056 [6, 8] and NGC 1068 [7] to constrain the DM-neutrino scattering cross-section. In these analyses, the 90% confidence level (CL) upper limits on the cross-section were derived by adopting the corresponding 90% CL lower limits on the number of neutrino events observed by IceCube for each source. In this work, we instead perform a full statistical treatment to derive constraints on the DM–neutrino cross-section, both for individual sources and through a stacked analysis of all four sources.
The rest of this article is structured as follows: In section II we report the astrophysical neutrino data in terms of the number of events observed from each source. We then discuss the spike density models predicting the DM distribution around SMBHs in section III. In section IV, we discuss the neutrino flux attenuation due to DM-neutrino scattering as the neutrinos propagate from AGN to the earth. In section V, we provide the details of the statistical analysis. We then present our model independent results in section V, followed by a model interpretation in section VI. Finally, we conclude in section VII.
II High Energy Neutrinos from Astrophysical Sources
[…]
$\Phi_{\nu_{\mu}+\bar{\nu}_{\mu}}=\Phi=\Phi_{ref}\bigg(\frac{E_{\nu}}{E_{0}}\bigg)^{-\Gamma},$ (1)
where $\Phi_{ref}$ is the normalization flux, $\Gamma$ is the spectral index, and $E_{\nu}$ and $E_{0}$ are the energy of neutrinos and the reference energy, respectively. Following this power law and the methodology of [23], we list the observed number of events $n_{s}$ for all four sources using the IceCube public data, in table 1.
[…]
Source $n_{s}\pm 1\sigma$ $\Gamma\pm 1\sigma$
NGC 1068 $75^{+16}_{-15}$ $2.75^{+0.2}_{-0.2}$
TXS 0506+056 $8^{+12}_{-12} 2.2$$2^{+0.3}_{-0.2}$
PKS 1424+240 $42^{+16}_{-14}$ $3.0^{+0.5}_{-0.4}$
NGC 4151 $30^{+13}_{-10}$ $2.4^{+0.4}_{-0.3}$
III DM Distribution Around SMBHs
In this section, we discuss the possibility of an enhanced DM distribution around SMBHs at the center of active galaxies. In early theoretical and empirical models, the DM density in the region close to SMBH was approximated as cuspy profile of the form [24, 25]
$\rho\simeq\rho_{s}\left(\frac{r}{r_{s}}\right)^{-\gamma},$ (2)
within the small range of radii $r$. Here, $\rho_{s}$ and $r_{s}$ are the characteristic density and scale radius, respectively and are defined as [24, 26, 27]
$\rho_{s}=\frac{\Delta}{3}\frac{\eta^{3}}{\ln(1+\eta)-\frac{\eta}{1+\eta}}\rho_{c};~~~~r_{s}=8.8\bigg(\frac{M_{vir}}{10^{11}M_{\odot}}\bigg)^{0.46}~{\rm kpc}.$ (3)
The quantity $\rho_{c}=9\times 10^{-27}$ kg/m3 is the critical density of the Universe, $\Delta$ is the virial over density $\approx 300$ [26], $M_{vir}$ is the virial mass, and the concentration parameter $\eta$ of the NFW halo is parameterized as
$\eta\simeq 13.6\bigg({\frac{M_{vir}}{10^{11}M_{\odot}}}\bigg)^{-0.13}.$
However, it has been reported that the possibility of adiabatic accretion of DM gives this profile a spiky nature that can be described as follows [28]
$\rho_{\chi}=\rho_{R}\bigg(1-\frac{4R_{s}}{r}\bigg)^{3}\bigg(\frac{R_{sp}}{r}\bigg)^{\alpha},$ (4)
where, $R_{s}=2GM_{BH}/c^{2}$ is the Schwarzschild radius, $R_{sp}$ is the DM spike radius, and $\rho_{R}\approx\rho(R_{sp})$ (using Eq. 2) represents the initial cuspy DM density profile at $r=R_{sp}$. The spectral index $\alpha$ is related to the parameter $\gamma$ of the initial cuspy profile as $\alpha=(9-2\gamma)/(4-\gamma)$. On the other hand, a cuspy DM density profile (the NFW-profile) [24] was established across a wide range of halo masses, ranging from dwarf galaxies to rich clusters, within the Cold Dark Matter (CDM) cosmological framework. This profile scales with halo radius (r) as
$\rho_{NFW}(r)=\rho_{s}\bigg(\frac{r}{r_{s}}\bigg)^{-\gamma}\bigg(1+\bigg(\frac{r}{r_{s}}\bigg)^{\xi}\bigg)^{-(\beta-\gamma)/\xi},$ (5)
where $(\xi,\beta,\gamma)=(1,3,1)$ [29, 30]. Here, $\gamma$ is the inner slope that controls the cusp/core behavior near the center, $\beta$ is the outer slope that determines the falloff at the radius $r$, and $\xi$ defines the sharpness of the transition between the inner and outer regimes. For the value of $\gamma=1$, as found in citation for the NFW profile, the resulting spike index is $\alpha=7/3$, which determines the spike density profile in Eq. 4.
Another possibility is gravitational scattering between the DM and dense stellar components, which is likely to be present in the inner region of the galaxy. This scenario relaxes the spike profile by lowering the spectral index to $\alpha=3/2$ independent of any value of $\gamma$ [31, 32, 33, 34].
A precise incorporation of these two possibilities [28, 32, 35] can be represented as
[…]
where $r_{i}=4R_{s}$ is the inner radius of the spike, and $r_{h}=GM_{BH}/\sigma^{\ast 2}$ is the influence radius of SMBH with $\sigma^{\ast}$ being the stellar velocity dispersion. The effects of DM scattering with stellar components are significant only within this radius [28]. $\rho_{N}$ and $\rho_{N}^{\prime}$ are normalization constants. Their evaluation is discussed later in this section. For $r>r_{h}$, the spike profile converges to
$\rho_{\chi}^{7/3}(r)\simeq\rho_{N}\!\left(1-\dfrac{4R_{S}}{r}\right)^{3}\!\left(\dfrac{R_{sp}}{r}\right)^{7/3},~~~~~r\geq r_{i}.$ (7)
Table 2: List of parameters used in this work associated with each source. We have also reported the $\Sigma_{\chi}$ values, in units of GeV/cm2, for each benchmark model from all sources for $m_{\chi}=1$ GeV and $r=10^{4}$ pc in this table.
Parameters NGC 1068 NGC 4151 TXS 0506+056 PKS 1424+240
$M_{BH}$ ($M_{\odot}$) $1.0\times 10^{7}$[36] $2.0\times 10^{7}$[37] $3.09\times 10^{8}$[38] $1.0\times 10^{9}$[39]
$t_{BH}$(yr) $10^{9}$ $10^{9}$ $10^{9}$ $10^{9}$
$R_{s}$(pc) $9.5\times 10^{-7}$ $1.9\times 10^{-6}$ $2.95\times 10^{-5}$ $9.5\times 10^{-5}$
$r_{h}$(pc) $6.5\times 10^{5}R_{s}$ $33.4\times 10^{5}R_{s}$ $10^{5}R_{s}$ $10^{5}R_{s}$
$R_{em}$(pc) $30R_{s}$ $30R_{s}$ $2.2\times 10^{3}R_{s}$ $6.7\times 10^{2}R_{s}$
$\Sigma_{\chi{\rm-BM1}}$ $5.77\times 10^{31}$ $2.88\times 10^{31}$ $2.69\times 10^{28}$ $3.95\times 10^{28}$
$\Sigma_{\chi{\rm-BM2}}$ $2.45\times 10^{28}$ $71\times 10^{28} 2.2$$27\times 10^{28}$ $3.15\times 10^{28}$
$\Sigma_{\chi{\rm-BM3}}$ $9.45\times 10^{26}$ $1.04\times 10^{27}$ $2.42\times 10^{27}$ $2.91\times 10^{27}$
$\Sigma_{\chi{\rm-BM1^{\prime}}}$ $9.28\times 10^{28}$ $4.64\times 10^{28}$ $6.04\times 10^{27}$ $3.49\times 10^{27}$
$\Sigma_{\chi{\rm-BM2^{\prime}}}$ $9.91\times 10^{27}$ $7.95\times 10^{27}$ $5.96\times 10^{27}$ $3.47\times 10^{27}$
$\Sigma_{\chi{\rm-BM3^{\prime}}}$ $1.11\times 10^{27}$ $1.0\times 10^{27}$ $2.39\times 10^{27}$ $1.78\times 10^{27}$
At sufficiently high DM densities, DM self-annihilations become significant. The process leads to a flattening of the DM spike towards the innermost region, in which the density reaches a saturation (maximum) value, $\rho_{sat}=m_{\chi}/(\langle\sigma_{a}v\rangle t_{BH})$. Here, $m_{\chi}$ is the DM mass, $\langle\sigma_{a}v\rangle$ is the thermally averaged annihilation cross section of DM, and $t_{BH}$ is the age of the SMBH. Taking these effects into account, the resulting DM spike density profile is expressed as [28]
[…]
where $\alpha=7/3,3/2$. In our analysis, we consider the benchmark models as given below [40]
  • BM1, BM2, BM3: choose $\alpha=7/3$ with $\langle\sigma_{a}v\rangle=(0,10^{-28},3\times 10^{-26})$ cm3 sec-1, respectively.
  • BM1, BM2, BM3: same annihilation cross-sections but with $\alpha=3/2$.
The choice $\langle\sigma_{a}v\rangle=0$ is applicable to scenarios such as asymmetric DM. In contrast, the value $\langle\sigma_{a}v\rangle=3\times 10^{-26}\text{ cm}^{3}\text{ sec}^{-1}$ corresponds to the canonical expectation for thermal relic DM.
III.1 Profile parameterization
Figure 1: DM column density $\Sigma$ Vs. $r_{l}$ (the lower limit to the integration given in Eq. (13)) for each source. We considered the upper limit to the same integration $r=10^{4}$ pc and $m_{\chi}=1$ GeV to obtain these plots. The shaded regions represent the possible range of $R_{\rm em}$ for each source.
The values of the parameters we have used in this work are provided in Table 2, for each source. For NGC 1068, where low-energy neutrinos are expected to originate primarily from hadronic interactions in the inner dusty torus or corona, we adopt an emission radius of $R_{em}=30R_{s}$, following [7, 41]. In the absence of direct observational constraints, a physically motivated range for these Seyfert systems spans a few tens of Schwarzschild radii [42, 13, 41, 43, 44, 45]. Owing to their comparable accretion structure and central engine geometry, we apply the same choice of $R_{em}$ to NGC 4151. The same argument also motivates using an identical $R_{em}$ for the blazars TXS 0506+056 [6] and PKS 1424+240, which are both gamma-ray blazars with similar black hole masses. Throughout our analysis, we assume a black-hole age of $10^{9}$ yr, a commonly adopted and well-supported timescale based on [46]. The parameter $r_{h}$ has been taken from [36, 47, 48, 49, 50] for NGC 1068, while for NGC 4151, we use the stellar dispersion velocity observed to be $\sigma^{\ast 2}=116\pm 3$ km/sec in [51] to calculate $r_{h}$. In the absence of any information on $r_{h}$ or $\sigma^{\ast}$ for TXS 0506+056 and PKS 1424+240, we have assumed $r_{h}=10^{5}R_{s}$, the typical radius relevant for BH mass estimations [52]. To calculate the normalizations $\rho_{N}$ and $\rho_{N}^{\prime}$ we need the mass of the BH, which is given as
$M_{BH}\approx 4\pi\int_{r_{i}}^{r_{h}}drr^{2}\rho_{\chi}(r).$ (9)
Here, $r_{h}$ is the influence radius of the BH, while the DM density $\rho_{\chi}(r)\simeq\rho_{\chi}^{\alpha}$ ($\alpha=7/3,3/2$) for $r~\gg~r_{i}$, given by Eqs. 6 and 7. To calculate $\rho_{N}$ we use the relation $\mathcal{M}\approx\rho_{N}r_{h}^{3/2}$ or $\mathcal{M}\approx\rho_{N}R_{sp}^{7/3}$. Under the condition $\rho_{\chi}(r)\simeq\rho_{\chi}^{\alpha}$ one can write
$\mathcal{M}\simeq\frac{M_{BH}}{4\pi\big(f_{\alpha}(r_{h})-f_{\alpha}(r_{i})\big)},$ (10)
where,
$f_{\alpha}(r)\equiv r^{-\alpha}\bigg[\frac{r^{3}}{3-\alpha}+\frac{12R_{s}r^{2}}{\alpha-2}-\frac{48R_{s}^{2}r}{\alpha-1}+\frac{64R_{s}^{3}}{\alpha}\bigg].$ (11)
This expression follows from integrating the binomial expansion of the factor
$\rho(r)\propto r^{2-\alpha}\left(1-\frac{4R_{S}}{r}\right)^{3},$
where the term $\left(1-4R_{S}/r\right)^{3}$ encodes the relativistic suppression of phase-space orbits near the BH. The leading contribution, proportional to $r^{2-\alpha}$, corresponds to the standard Newtonian adiabatic spike term, while the remaining terms scaling as $r^{1-\alpha}$, $r^{-\alpha}$ and $r^{-1-\alpha}$ arise from the linear, quadratic, and cubic components of the expansion of $(1-4R_{S}/r)^{3}$, and therefore represent the first, second, and third order general relativistic (GR) corrections to the spike structure respectively [28, 53].
However, we do not consider the relativistic corrections to the spike as proposed in [53] because they are negligible at distances $\gtrsim 20$ pc [7]. Indeed for all sources examined here, the value of $R_{em}$ exceeds this limit. $\rho_{N}^{\prime}$ is calculated using the relation $\rho_{N}^{\prime}\simeq\rho_{N}(r_{h}/R_{sp})^{7/3}$ ensuring that the profile $\rho_{\chi}^{3/2}$ matches at the intersection point $r_{h}$. In addition, the spike radius can be calculated using the condition that the DM halo should continue following the NFW profile outside of the spike radius by imposing the condition $\rho_{N}^{\prime},\rho_{N}\simeq\rho_{s}(R_{sp}/r_{s})^{-\gamma}$. Hence, we define the spike radii as $R_{sp}^{3/2}=(\mathcal{M}/\rho_{s}r_{s})^{3/4}~r_{h}^{5/8}$ and $R_{sp}^{7/3}=(\mathcal{M}/\rho_{s}r_{s})^{3/4}$.
The parameters $r_{s}$ and $\rho_{s}$ defining the NFW profile, are calculated based on the virial mass, which is taken as the DM halo mass $M_{DM}$ and is calculated using the relation of $M_{DM}$ with the SMBH mass as
$M_{BH}\sim 7\times 10^{7}M_{\odot}\bigg(\frac{M_{DM}}{10^{12}M_{\odot}}\bigg)^{4/3},$ (12)
[…]
The emission point of neutrinos $R_{em}$ near the SMBH is taken as $30\,R_{s}$ for NGC 1068 and NGC 4151 because of lower-energy (<10<10 TeV) neutrinos detected from these sources. This requires relatively lower-energy protons to interact with a higher-energy radiation field or dense gas, both of which are found relatively closer to the SMBH, (see e.g., [42]). While for TXS 0506+056, neutrinos having higher energy (290 TeV) are supposedly produced at a comparatively larger distance from the SMBH, where the radiation field is at relatively lower energy, thus interacting with high-energy protons, which is consistent with the lepto-hadronic model of high energy neutrino production in blazars [15].
III.2 DM column density
Combining the ingredients described in the previous sections, we evaluate the DM column density, which quantifies the probability for neutrinos to scatter off DM as they propagate towards the Earth. The column density $\Sigma_{\chi}$ is obtained by integrating $\rho_{\chi}$ over an estimated distance traveled by neutrinos, expressed as
$\Sigma_{\chi}(r)=\int_{r_{l}}^{r}dr^{\prime}\rho_{\chi}(r^{\prime})\,.$ (13)
Here, $r_{l}=R_{em}$ with $r=10^{4}$ pc the upper limit of the integral given in Eq. (13). Above this upper limit, we obtain an approximately constant $\Sigma_{\chi}$ for each source. Fig (1) shows the variation of column density $\Sigma_{\chi}$ Vs. $r_{l}$ for each source.
In this analysis, a key consideration is the uncertainty associated with the neutrino emission point $R_{\rm em}$. We selected specific values for $R_{\rm em}$ (see Table 2) based on previous analyses [6, 7]. In Fig. 1, we observe how $\Sigma_{\chi}$ varies with different values of $R_{\rm em}$. The shaded blue region represents the allowed range of $R_{\rm em}$ for each source. For instance, for NGC 1068 and NGC 4151, the range is $10R_{s}\lesssim R_{\rm em}\lesssim 90R_{s}$ [45]. In contrast, for TXS 0506+056 and PKS 1424+240, we considered $R_{\rm em}$ in the range of $(10^{-3}-10^{-1})$ pc as indicated in Ref. [38]. The constraints for NGC 1068 and NGC 4151 can vary significantly within the specified range of $R_{\rm em}$ for the BM1 profile, potentially by an order of magnitude. A minor variation is observed in $\Sigma_{\chi}$ for BM1; however, $\Sigma_{\chi}$ for other profiles remain almost constant within the considered $R_{\rm em}$ range. For TXS 0506+056 and PKS 1424+240, the BM1 spike may vary within nearly three orders of magnitude, while the BM1 shows a decrease of about one order of magnitude. Additionally, slight variations are noted in BM2 and BM2, whereas the spikes for BM3 and BM3 are almost constant. Since the bound on $\sigma_{0}$ is inversely proportional to the column density $\Sigma_{\chi}$, a decrease in $\Sigma_{\chi}$ will relax the bound on $\sigma_{0}$ by the same order of magnitude.
IV Neutrino flux attenuation due to interaction with DM
The neutrino flux attenuation due to scattering with DM is evaluated using the cascade equation [3, 57]
$\frac{d\Phi}{d\tau}(E_{\nu})=-\,\sigma_{\nu\chi}\,\Phi(E_{\nu})+\int_{E_{\nu}}^{\infty}dE^{\prime}_{\nu}\,\frac{d\sigma_{\nu\chi}}{dE_{\nu}}(E^{\prime}_{\nu}\!\to\!E_{\nu})\,\Phi(E^{\prime}_{\nu})\,$, (14)
where, $\tau=\Sigma_{\chi}(r)/m_{\chi}$ is the accumulated column density over the DM mass, $E_{\nu}$ and $E_{\nu}^{\prime}$ are the outgoing and incoming neutrino energies, respectively. The first term in this equation represents the attenuation of the neutrino flux, while the second term represents the redistribution of high energy neutrinos in the lower energy bins, due to scattering with DM.
Figure 2: Constraints on $\sigma_{0}$ for the constant $\sigma_{\nu{\chi}}$ for all the benchmark models for all the sources. The blue-dotdashed line represents the bound from the stacking analysis combining the data from all four sources. The constraints obtained by Cline et al. for NGC 1068 [7] (black dashed) and TXS 0506+056 [6] (black dotted) are shown for comparisons. For context, we also overlay the model independent constraints on the DM-neutrino scattering cross-section coming from SN 1987A [58] (magenta), stellar-neutrinos [59] (gray solid), diffuse supernova (dSN) Xenon1T (yellow) and dSN Super-K (purple solid) [60], CMB+BAO [61] (green dotted), Lyman-$\alpha$ [5] (orange dotted).
Figure 3: Constraints on $\sigma_{0}$ for the energy-dependent $\sigma_{\nu\chi}$ for all the benchmark models for all the sources. The solid-blue line represents the bound from the stacking analysis combining the data from all four sources. The limits obtained previously [7, 6] have been rescaled to $E_{0}=10$ TeV for comparisons.
In this work, we consider two scenarios: (i) constant $\sigma_{\nu\chi}=\sigma_{0}$, and (ii) linearly energy dependent $\sigma_{\nu\chi}=\sigma_{0}\frac{E_{\nu}}{E_{0}}$. In the case of constant cross-section, the second term is zero, hence, the attenuated flux is $\Phi=\Phi_{\nu+\bar{\nu}}~e^{-\sigma_{0}\Sigma_{\chi}/m_{\chi}}$. However, in case of energy-dependent $\sigma_{\nu\chi}$, the energy re-distribution term cannot be neglected, and an analytic solution to the cascade equation above is highly non-trivial. Following [6], we use a discretized form of this equation for each energy bin $i$ given by
$\frac{d\Phi_{i}}{dy}=A\!\left[-\,\hat{E}_{i}\,\Phi_{i}+\Delta x\,\ln 10\sum_{j=i}^{N}\hat{E}_{j}\,\Phi_{j}\right],$ (15)
where, $y$ is a unitless quantity that varies between 0 and $1$, $\hat{E}_{i}=10^{x_{i}}$, $x=\log_{10}(E_{\nu}/TeV)$ and
$A=\frac{\Sigma_{\chi}}{m_{\chi}}\,\frac{\sigma_{0}}{\hat{E}_{0}}.$ (16)
For each source, we adopt a reference scale energy $E_{0}=10$ TeV, which provides a consistent basis for comparing all sources and enables a stacking analysis that combines data across the full sample.
V Data analysis and Results
V.1 Likelihood-based $\chi^{2}$ analysis
To place constraints on the DM-neutrino scattering cross-section, from individual sources, we define a likelihood-based $\chi^{2}$ using Poisson statistics as follows [62]
$\chi^{2}=2\sum_{i}\bigg[N_{i}^{th}-N_{i}^{ex}+N_{i}^{ex}\ln\bigg(\frac{N_{i}^{ex}}{N_{i}^{th}}\bigg)\bigg]+2\sum_{j\neq i}N_{j}^{th}$ (17)
where $i$ corresponds to the energy bins for which the experimental event number $N_{i}^{ex}\neq 0$ and $j$ corresponds to the energy bins for which $N_{j}^{ex}=0$. The quantities $N_{i}^{th}$ or $N_{j}^{th}$ correspond to the theoretical estimate of the number of neutrino events from the flux model, after taking into account DM-neutrino scattering and standard neutrino oscillations over astrophysical distances. Here $N^{ex}=n_{s}$ as listed in Table 1. We perform a stacking analysis by combining $\chi^{2}$ from all sources (NGC 1068, TXS 0506+056, PKS 1424+240 and NGC 4151), i.e.
$\chi^{2}_{\rm total}=$ $\chi^{2}_{\rm(NGC~1068)}+\chi^{2}_{\rm(TXS~0506+056)}$
$+\chi^{2}_{\rm(PKS~1424+240)}+\chi^{2}_{\rm(NGC~4151)},$
which we minimize for a common $\sigma_{0}$ and $m_{\chi}$.
V.2 Results for constant cross-section
The results of this analysis for constant $\sigma_{\nu\chi}$ are shown in Fig. 2.
  • We find that for benchmark models BM1 and BM1, for which $\langle\sigma_{a}v\rangle=0$ and consequently the column density $\Sigma_{\chi}(m_{\chi})$ is nearly independent of the DM mass $m_{\chi}$, the strongest constraint arises from the NGC 1068 source. To understand the reason behind this result, we evaluate the $\chi^{2}$ given in Eq. (17) for the constant cross-section as follows
    $\chi^{2}\approx n_{s}\big[e^{-\sigma_{0}\Sigma_{\chi}/m_{\chi}}-1+\sigma_{0}\Sigma_{\chi}/m_{\chi}\big],$ (19)
    where $n_{s}$ is the total number of neutrino events over all energy bins. From Eq. (19), the 90% upper bound on the DM-neutrino cross-section is obtained as
    $\sigma_{0}\lesssim cm_{\chi},$
    where, $c\propto\frac{1}{\sqrt{n_{s}}\Sigma_{\chi}}$. The strongest bound observed for NGC 1068 is attributed to the fact that the factor $c$ calculated from the $\chi^{2}$ analysis is the smallest for NGC 1068, due to the observed number of events for this source being the highest. We also find that for more massive SMBHs, the DM density $\rho_{\chi}$ peaks at larger radii, with lower spike height, leading to a smaller column density $\Sigma_{\chi}$ along the neutrino’s propagation to the Earth. Since NGC 1068 hosts the least massive SMBH among the four sources considered, it correspondingly yields the largest $\Sigma_{\chi}$. Consequently, a combination of the largest signal $n_{s}$ and the largest column density produces the strongest constraints for this source. As a result, NGC 1068 largely dominates the results of the stacking analysis.
  • For the benchmark models BM2, BM3, BM2 and BM3, where DM-annihilation effects are non-zero, PKS 1424+240 provides the strongest constraint for most of the range of $m_{\chi}$, as shown in Fig. 2. The reason behind this is that DM spikes with higher density get depleted faster due to DM self-annihilation resulting in large decrease in $\Sigma_{\chi}$ as shown for NGC 1068 and NGC 4151 in Fig. 1 and discussed in section III.2. In comparison, the decrease in $\Sigma_{\chi}$ for TXS 0506+056 and PKS 1424+240 at the emission radii of neutrinos (blue shaded region in Fig. 1) is smaller for most $m_{\chi}$ values because of their more massive black holes. Because of a larger $\Sigma_{\chi}$ compared to the NGCs, PKS 1424+240 and TXS 0506+056 give stronger constraints on $\sigma_{0}$ in these benchmark models for most of the $m_{\chi}$ range.
V.3 Results for energy dependent cross-section
In this section, we perform a $\chi^{2}$ analysis for a linearly energy-dependent scattering cross-section, of the form $\sigma_{\nu\chi}=\sigma_{0}\frac{E_{\nu}}{E_{0}}$. We observe the following points:
  • For benchmark models BM1 and BM1 (with $\langle\sigma_{a}v\rangle=0$), the strongest constraint is again obtained from NGC 1068, yielding an upper limit of roughly $\sigma_{0}\leq 10^{-39}$ cm2. This bound is slightly more stringent than that obtained in the energy-independent (constant cross-section) scenario.
  • For models BM2, BM3, BM2 and BM3 (with $\langle\sigma_{a}v\rangle\neq 0$), the strongest constraint is obtained from TXS 0506+056, in contrast to the constant cross-section case where the strongest bound comes from PKS 1424+240. The difference is attributed to the energy dependence of the interaction which makes the neutrino energy event distribution particularly important. Since the event distribution for TXS 0506+056 peaks at higher energy, compared to that of PKS 1424+240, it yields the most stringent constraint in the energy-dependent scenario. The event distributions for all sources are shown in Fig. 6 in the Appendix.
  • In reference [6], the results are obtained for TXS 0506+056 considering the neutrino flux predicted by the lepto-hadronic model studied in [63], which estimated the number of events to be approximately 2. The 90% CL lower limit is set to be $\sim$ 0.1 at a fixed reference energy of $E_{0}=290$ TeV. In our analysis, we perform a full statistical treatment by assuming that the neutrino flux follows the unbroken power law given in Eq. (1), to derive the lower bound on the number of events over a broad energy range from 0.1 TeV - 1 PeV for each source. With this approach, our results for TXS 0506+056 are in very good agreement with those of reported in [6].
VI Model Interpretation
In this section, we will interpret our results in the parameter space of a well-motivated model of DM with significant coupling to neutrinos. We consider a scenario in which the SM is extended to a gauged and anomaly-free $U(1)_{L_{\mu}-L_{\tau}}$ group [64, 65]. Spontaneous symmetry breaking leads to a massive gauge boson $Z^{\prime}$ that couples primarily to $2^{\rm nd}$ and $3^{\rm rd}$ generation SM leptons ($\mu$, $\tau$, $\nu_{\mu}$, $\nu_{\tau}$) and their anti-particles.
The Lagrangian of interactions is given by
$\mathcal{L}\supset\frac{m_{Z^{\prime}}^{2}}{2}Z_{\mu}^{\prime}Z^{\prime\mu}+Z_{\mu}^{\prime}(g_{f}\mathcal{J}_{f}^{\mu}+\varepsilon e\mathcal{J}_{EM}^{\mu}+g_{\chi}\mathcal{J}_{\chi}^{\mu}),$ (20)
where $m_{Z^{\prime}}$ denotes the gauge boson mass, and $g_{f}\equiv Q_{\mu-\tau}g_{\mu-\tau}$ is the gauge coupling to the SM $2^{\rm nd}$ and $3^{\rm rd}$ generation leptons. Throughout this work, we assume a unit charge, $Q_{\mu-\tau}=1$. The $L_{\mu}-L_{\tau}$ current is given by
$\mathcal{J}_{f}^{\mu}$ $=\bar{\mu}\gamma^{\mu}\mu+\bar{\nu}_{\mu}\gamma^{\mu}P_{L}\nu_{\mu}-\bar{\tau}\gamma^{\mu}\tau-\bar{\nu}_{\tau}\gamma^{\mu}P_{L}\nu_{\tau},$ (21)
where $P_{L}=\frac{1}{2}(1-\gamma^{5})$ is the left handed chirality operator. The kinetic mixing $\varepsilon$, between $Z^{\prime}$ and SM gauge bosons is generated at the 1-loop level and is given by [66, 67]
$\varepsilon$ $=-\frac{eg_{\mu-\tau}}{12\pi^{2}}{\rm log}\left(\frac{m_{\tau}^{2}}{m_{\mu}^{2}}\right)$
$\simeq-\frac{g_{\mu-\tau}}{70},$ (22)
with the electromagnetic current for all SM fermions represented by $\mathcal{J}_{EM}^{\mu}=\bar{f}\gamma^{\mu}f$. Although kinetic mixing induces a coupling of all other SM fermions to the $Z^{\prime}$, these interactions are subdominant. We consider a DM $\chi$ charged under this symmetry, with $g_{\chi}\equiv Q_{\mu-\tau}^{\chi}g_{\mu-\tau}$ is the coupling of $Z^{\prime}$ to DM. We assume that DM is vector-like under this gauge symmetry such that the charge ($Q_{\mu-\tau}^{\chi}$) can differ from unity [68, 69]. For illustration, we focus on two benchmark DM scenarios (pseudo-Dirac fermion and complex scalar), in which the corresponding current is given by
$\mathcal{J}_{\chi}^{\mu}=\left\{\begin{array}[]{lc}\bar{\chi}_{1}\gamma^{\mu}\chi_{2}&{\rm Pseudo-Dirac}\\ i\chi^{*}\partial^{\mu}\chi+c.c&{\rm Complex~Scalar}.\end{array}\right$. (23)
Figure 4: Limits on the model parameters for energy-dependent scattering cross sections are shown for pseudo-Dirac dark matter (left panel) and complex scalar dark matter (right panel). In each panel, the black solid and dot-dashed curves indicate parameter values that reproduce the observed dark matter relic abundance with $\Omega h^{2}=0.12$ [70]. The colored curves show the constraints obtained from our stacking analysis for the different SMBH spike density benchmark models considered. The gray shaded regions denote existing bounds on the parameter space from complementary probes, including constraints from dark matter annihilation into neutrinos [69], as well as limits from early-Universe cosmology, astrophysical observations, and terrestrial accelerator searches for new gauge bosons in this class of models [67].
Figure 5: Limits on the dark matter-neutrino scattering cross-section vs dark matter mass. We include the freeze-out relic density in our two dark matter scenarios of interest. We overlay our stacking analysis results for each benchmark model spike density model.
VI.1 Dark Matter Relic Density
We consider thermal relic DM whose present-day abundance is set by the freeze-out mechanism. We focus on the regime $m_{Z^{\prime}}>m_{\chi}$, with $m_{\chi}$ the DM mass, such that the dominant annihilation channel is $\bar{\chi}\chi\rightarrow\bar{f}f$. In both model scenarios considered above, the corresponding annihilation cross-section (in terms of Mandelstam s), for $\chi$ into $f=\mu,\tau$, is given by
$\sigma_{ann}(s)=\frac{g_{\chi}^{2}g_{\mu-\tau}^{2}}{12\pi s}\frac{\sqrt{s-4m_{f}^{2}}(s+2m_{f}^{2})(s+2m_{\chi}^{2})}{(s-4m_{\chi}^{2})[(s-m_{Z^{\prime}}^{2})^{2}+m_{Z^{\prime}}^{2}\Gamma_{Z^{\prime}}^{2}]},$ (24)
where $m_{f}$ is the SM fermion mass, and the total $Z^{\prime}$ decay width to $\mu$ and $\tau$ is given by
$\Gamma_{Z^{\prime}}=\frac{g_{\mu-\tau}^{2}m_{Z^{\prime}}}{12\pi}\left(1+\frac{2m_{f}^{2}}{m_{Z^{\prime}}^{2}}\right)\sqrt{1-\frac{4m_{f}^{2}}{m_{Z^{\prime}}^{2}}}.$ (25)
The decay width into neutrinos is $\Gamma_{Z^{\prime}}\simeq\frac{g_{\mu-\tau}^{2}m_{Z^{\prime}}}{24\pi}$. In the non-relativistic limit, the thermally averaged annihilation cross-section can be parametrized as $\langle\sigma_{a}v\rangle\equiv\sigma_{a}\,x^{-n}$[71, 68, 69], where $x\equiv\frac{m_{\chi}}{T}$ is a parameter representing the temperature. The parameter n characterizes the dominant partial wave of the annihilation process: $n=0$ for s-wave annihilation (as in the pseudo-Dirac scenario) and $n=1$ for p-wave annihilation (as in the complex scalar case). Hence, the thermally averaged annihilation cross-section for both scenarios is generalized as
$\sigma_{a}^{f}=\frac{g_{\mu-\tau}^{2}g_{\chi}^{2}}{k\pi m_{\chi}^{2}}\frac{(2+m_{f}^{2}/m_{\chi}^{2})\sqrt{1-m_{f}^{2}/m_{\chi}^{2}}}{[(4-m_{Z^{\prime}}^{2}/m_{\chi}^{2})^{2}+m_{Z^{\prime}}^{2}\Gamma_{Z^{\prime}}^{2}/m_{\chi}^{4}]},$ (26)
and into neutrinos;
$\sigma_{a}^{\nu}=\frac{g_{\mu-\tau}^{2}g_{\chi}^{2}}{k\pi m_{\chi}^{2}}\frac{1}{[(4-m_{Z^{\prime}}^{2}/m_{\chi}^{2})^{2}+m_{Z^{\prime}}^{2}\Gamma_{Z^{\prime}}^{2}/m_{\chi}^{4}]},$ (27)
with $\sigma_{a}=\sigma_{a}^{f}+\sigma_{a}^{\nu}$. We follow the freeze-out formalism in [69] and obtain the analytic approximation of the relic abundance of $\chi$ as
$\Omega_{\chi}h^{2}=8.77\times 10^{-11}\frac{l(n+1)x_{f}^{n+1}{\rm GeV^{-2}}}{(g_{*,s}/\sqrt{g_{*}})\sigma_{a}},$ (28)
where $l=2$ if DM has an antiparticle and $l=1$ if it is its own antiparticle. $g_{*,s}$ and $g_{*}$ are the entropic degrees of freedom and the relativistic degrees of freedom, respectively, with $x_{f}$ the freeze-out temperature. For the models described above, the factors $(n,k,l)=(0,2,2)$ and $(1,12,2)$ for the pseudo-dirac and complex scalar models, respectively  [^1] Adopting the precise treatment of [72], for our benchmark parameters, has a negligible impact on our results.
VI.2 Dark matter - Neutrino Scattering
Neutrinos with energy $E_{\nu}$ can scatter off ambient DM, which we take to be at rest in the Galactic frame. These interactions can modify the propagation of high-energy neutrinos leading to potentially observable effects in neutrino telescopes. Within the DM scenarios introduced above, the DM-neutrino scattering proceeds via exchange of the new gauge boson $Z^{\prime}$. At leading order, the elastic scattering cross-section for pseudo-Dirac DM is given by
$\sigma_{\chi\nu}=\frac{g_{\mu-\tau}^{2}g_{\chi}^{2}}{32\pi E_{\nu}^{2}m_{\chi}^{2}}\Bigg[(m_{Z^{\prime}}^{2}+m_{\chi}^{2}+2E_{\nu}m_{\chi})\,{\rm log}\left(\frac{m_{Z^{\prime}}^{2}(2E_{\nu}+m_{\chi})}{m_{\chi}(4E_{\nu}^{2}+m_{Z^{\prime}}^{2})+2E_{\nu}m_{Z^{\prime}}^{2}}\right)$ (29)
$+4E_{\nu}^{2}\left(1+\frac{m_{\chi}^{2}}{m_{Z^{\prime}}^{2}}-\frac{2E_{\nu}(4E_{\nu}^{2}m_{\chi}+E_{\nu}(m_{\chi}^{2}+2m_{Z^{\prime}}^{2})+m_{\chi}m_{Z^{\prime}}^{2})}{(2E_{\nu}+m_{\chi})(m_{\chi}(4E_{\nu}^{2}+m_{Z^{\prime}}^{2})+2E_{\nu}m_{Z^{\prime}}^{2})}\right)\Bigg].$
For complex scalar DM, the cross-section is given by
$\sigma_{\chi\nu}=\frac{g_{\mu-\tau}^{2}g_{\chi}^{2}}{16\pi E_{\nu}^{2}m_{\chi}}\left[\frac{4E_{\nu}^{2}m_{\chi}}{m_{Z^{\prime}}^{2}}+(m_{\chi}+2E_{\nu})\,{\rm log}\left(\frac{m_{Z^{\prime}}^{2}(2E_{\nu}+m_{\chi})}{m_{\chi}(4E_{\nu}^{2}+m_{Z^{\prime}}^{2})+2E_{\nu}m_{Z^{\prime}}^{2}}\right)\right].$ (30)
Both cross-sections were obtained following the description in [3] and capture the full energy dependence of high energy neutrinos. However, they exhibit limiting behaviors, which are expressed as
$\sigma_{\chi\nu}=\frac{g_{\mu-\tau}^{2}g_{\chi}^{2}}{\kappa\pi\,m_{Z^{\prime}}^{2}}\left\{\begin{array}[]{lc}1&E_{\nu}\gg m_{Z^{\prime}}/m_{\chi}\\ m_{\chi}E_{\nu}/m_{Z^{\prime}}^{2}&E_{\nu}\ll m_{Z^{\prime}}/m_{\chi},\end{array}\right$. (33)
where $\kappa=8$ for the pseudo-dirac DM and $\kappa=4$ for the complex scalar DM  [^2] We have verified that our calculations match the results of [3] up to a prefactor of $\frac{1}{2}$, which comes from the $P_{L}$ operator in the $Z^{\prime}$-$\nu$ coupling.. Equation (33) justifies our choices of $\sigma_{\chi\nu}=\sigma_{0}$ or $\sigma_{\chi\nu}\sim\sigma_{0}\,E_{\nu}$, as assumed throughout the study and illustrated in Figs. 2 and 3.
In Fig. 4, we show the results of our stacking analysis overlaid on the $U(1)_{L_{\mu}-L_{\tau}}$ model parameter space, plotting the gauge coupling $g_{\mu-\tau}$ vs gauge boson mass $m_{Z^{\prime}}$. On the left and right panels we showcase the results of the pseudo-Dirac fermion and complex scalar DM, respectively, assuming $m_{Z^{\prime}}=3m_{\chi}$ and $g_{\chi}=1$. The color lines represent our constraints incorporating the different spike density profiles after stacking the results from TXS 0506+056, PKS 1424+240, NGC 4151 and NGC 1068 (the blue dot-dashed line in each panel of Figs. 2 and 3). The solid blue and red lines represent the BM1 and $\rm BM1^{\prime}$ models, respectively. The dashed lines correspond to the BM2 and $\rm BM2^{\prime}$ models, while the dotted lines represent the BM3 and $\rm BM3^{\prime}$ spike density models, respectively.
The gray shaded regions indicate constraints on the model arising from various experimental and observational searches. The regions associated with the neutrino experiments (i.e. Borexino, KamLand, Super-K) correspond to parameter space in which galactic center DM annihilates into neutrinos; see [69] for further details. Constraints from early-universe cosmology, astrophysical observations and terrestrial probes, including the latest results from the muon g-2 measurements, are discussed in [67].
In this model, DM can also interact with electrons or nucleons, via kinetic mixing. However, constraints from direct detection experiments on this space are relatively weak (due to the suppressed kinetic mixing) and are therefore not shown. For brevity, we also omit other existing model constraints that are subdominant compared to those presented here.
Finally, in the left panel, the black solid line denotes the DM relic density predicted by the pseudo-Dirac fermion model. In the right panel, the dot-dashed line shows the dark matter relic abundance in the complex scalar model. Within the parameter space shown, the IceCube stacking analysis assuming the BM1 spike density profile yields the most stringent constraints on the gauge coupling. In particular, it constrains a substantial portion of the relic density in the pseudo-Dirac fermion scenario and effectively the entirety of the relic density-favored region in the complex scalar case.
Fig. 5 illustrates the impact of our stacking analysis on the DM parameter space, compared to other model-independent bounds by recasting the thermal relic density in the $\sigma_{0}$ vs $m_{\chi}$ plane. The black solid and dot-dashed curves correspond to the pseudo-Dirac and complex scalar models respectively, with $m_{Z^{\prime}}=3m_{\chi}$ and $g_{\chi}=1$. For illustration, we present the results in the limit $E_{\nu}\ll m_{Z^{\prime}}/m_{\chi}$ corresponding to the energy-dependent scattering cross-section scenario shown in Fig. 3. The color shaded regions indicate the model-independent bounds as discussed above. Overall, we find that our particle physics model-independent stacking analysis provides stringent constraints on the DM parameter space for the chosen benchmark parameters.
VII Discussion & Conclusions
We analyzed high energy astrophysical neutrino data from AGN and blazars, as observed by IceCube, to place constraints on the DM-neutrino scattering cross-section. In contrast to the current literature, where individual sources were analyzed and constraints derived from these, we perform a statistically robust stacking analysis in which we combine the results from 4 individual sources of extragalactic high energy neutrinos, namely TXS 0506+056, PKS 1424+240, NGC 4151 and NGC 1068.
Consistent with the literature, we perform our analysis for both a constant cross-section and a cross-section linearly dependent on the neutrino energy. The strongest constraints observed in our analysis are $\sigma_{0}\lesssim 8\times 10^{-39}$ cm2 and $\sigma_{0}\lesssim 10^{-39}$ cm2 in case of constant and energy-dependent cross sections, respectively, at 90% CL for 1 keV DM mass, which is derived from the stacking analysis for BM1 spike density profile.
For context, it is useful to compare our results with existing bounds in the literature. Previous cosmological analyses based on Milky Way satellite galaxy counts have placed upper limits of $\sigma_{\rm DM-\nu}\lesssim 4\times 10^{-34}\,{\rm cm}^{2}(m_{\rm DM}/{\rm GeV})$ at 95% CL for a constant DM–neutrino scattering cross section, corresponding to relic neutrinos with characteristic energies of order $10^{-4}\,\mathrm{eV}$ [82]. These constraints arise from the impact of DM–neutrino interactions on small-scale structure formation in the early Universe. In contrast, our analysis probes the same interaction using high-energy astrophysical neutrinos detected by IceCube, extending up to TeV energies, and provides the most stringent limits to date in this high-energy regime. For completeness, we also note that a comparatively weaker bound, of order $10^{-22}\,{\rm cm}^{2}\,{\rm GeV}^{-1}$, has been reported using the highest-energy event observed by KM3NeT, where only propagation through the Milky Way dark matter halo was considered [83].
Our results from the stacking analysis, in case of constant cross-section, are dominated by NGC 1068 due to the largest observed number of events and the expected large column density for this source. The benchmark model BM1, however, does not take into account effects such as gravitational scattering of DM with surrounding stars or self-annihilation of DM. These effects have been incorporated in other benchmark models BM1, BM2/3 and so on, and we found that our constraints on the DM-neutrino scattering cross-section are weaker for those models. Additionally, for those benchmark models, we found that the most stringent constraints on the DM-neutrino scattering cross-section come from sources with heavier SMBHs, making TXS 0506+056 and PKS 1424+240 particularly significant in this context.
In the energy-dependent cross-section scenario, the most stringent constraint arises for spike profile BM1 from NGC 1068, which is stronger than that derived in the energy-independent (constant cross-section) case. In contrast, for DM density profiles surrounding SMBHs with a nonzero $\langle\sigma_{a}v\rangle$ ($i.e.$, BM2, $\rm BM2^{\prime}$, BM3, $\rm BM3^{\prime}$), the strongest constraint instead comes from TXS 0506+056. This can be attributed to the detection of the highest-energy neutrinos from this source (with energies up to $\approx$290 TeV), which enhances experimental sensitivity in the energy-dependent regime.
In our analysis of model interpretations, we discussed the anomaly-free $U(1)_{L_{\mu}-L_{\tau}}$ model and presented our results from a stacking analysis conducted in the model’s parameter space. This model considers the scenario of pseudo-Dirac and complex scalar DM, where we established the most stringent constraints on the coupling and mediator mass parameters. A recent analysis of DM–neutrino interactions has examined scalar and fermionic DM coupled through various mediators, considering laboratory, cosmological, and astrophysical constraints, where they excluded many previously accepted scenarios [84]. On the other hand, our study focuses on a well-motivated anomaly-free model, deriving high-energy constraints in its parameter space from IceCube observations at TeV energies.
In conclusion, the constraints obtained in this study on DM-neutrino scattering cross-sections from individual sources are significantly enhanced when we conduct a stacking analysis for each assumed DM density profile around SMBHs. Collectively, these findings provide the most stringent constraints on the DM-neutrino scattering cross-section in this context to date.
Acknowledgements.
We thank Aaron Vincent, Gabrijela Zaharijas and Matteo Puel for helpful discussions. K.D. and S.R. were partially supported by a National Research Foundation (NRF) of South Africa grant facilitated through the National Institute of Theoretical and Computational Sciences (NITheCS). GM is supported by the Natural Sciences and Engineering Research Council of Canada (NSERC). During the completion of this work, GM was also supported by the University of California Irvine, School of Physical Sciences visiting fellowship as well as the Aspen Center for Physics, which is supported by the National Science Foundation grant PHY-2210452.
Appendix A Appendixes
Here we show the event distributions in the energy-dependent cross-section case (see Eq. (15)) for the four sources we considered in our analysis. The values for the constant $A$ in Eq. (15) for these plots are 0.0091 (NGC 1068), 0.0114 (NGC 4151), 0.0056 (TXS 0506+056) and 0.0311 (PKS 1424+240). These results are obtained keeping the spectral index $\Gamma$ fixed to the values derived from the IceCube public data for each source, as listed in Table 1. However, additionally, we found that adjusting the profiling over $\Gamma$ does not significantly impact our results.
Figure 6: Event distributions obtained from the IceCube data in the standard scenario (blue-solid) and in the DM-neutrino interaction scenario with energy-dependent cross-section at 90% CL (red-dotdashed)